OpenStudy (anonymous):

find the value of k so that the line has slope m: 6x+ky = 10 and m=-2 k=-8 k=4 k=3 k=-3

OpenStudy (anonymous):

ky = -6x + 10 -> y = (-6/k)x + 10/k So, -6/k = -2 -> k = 3

OpenStudy (anonymous):

thanks! what about (k+1)x+2y=6 and m=-(1/3)

OpenStudy (anonymous):

2y = -(k + 1)x + 6 -> y = -[(k + 1)/2]x + 3 -[(k + 1)/2] = -(1/3) -> [(k + 1)/2] = (1/3) k + 1 = 2/3 -> k = -1/3

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

You're welcome!